OpenStudy (anonymous):

Algebra 1, help please! (:

OpenStudy (anonymous):

Question: Explain how a four-term polynomial is factored by grouping and when a quadratic trinomial can be factored using this method. Include examples in your explanation.

OpenStudy (amoodarya):

\[2x^3-4x^2+2x-4 =\\2x^3+2x -4x^2-4 =\\2x(x^2+1) -4(x^2+1)=\\(x^2+1)(2x-4)\]

OpenStudy (anonymous):

Don't really understand what you just wrote.. o: Could you explain please? (:

OpenStudy (the_fizicx99):

He just grouped them

OpenStudy (anonymous):

Oh okay, thanks! (:

OpenStudy (the_fizicx99):

yw!